OpenStudy (anonymous):

calcII y=x^2, x=1, y=0 around x=2 using the shell method. ans 5pi/6 drawing below.

OpenStudy (anonymous):



OpenStudy (tkhunny):

What's your plan? \(2\pi\int\limits_{??}^{??} Radius\cdot Height\;dx\)

OpenStudy (anonymous):

pretty much

OpenStudy (tkhunny):

Okay, fill in the blanks.

OpenStudy (anonymous):

i've figured out most of the pieces.

OpenStudy (tkhunny):

What are the limits of the integration?

OpenStudy (anonymous):

2π∫(0,1)1⋅x^2dx

OpenStudy (tkhunny):

Not quite. Radius is (2-x)

OpenStudy (anonymous):

should it be in terms of y? that's what my book says.

OpenStudy (tkhunny):

Limits are also [1,2]. dy for shells? No.

OpenStudy (anonymous):

no, your misunderstanding the question.

OpenStudy (tkhunny):

Yup. I wraped the wrong piece. It's still not dy.

OpenStudy (anonymous):